The width of a rectangle is 3 ft less than the length. The area of the rectangle is 108ft^2. Find the length L and width W of the rectangle.
Answer by ikleyn(52778)

.

    W*(W+3) = 108  ft^2.


    W^2 + 3w - 108 = 0

Factor

    (W+12)*(W-9) = 0


W= 9 is the only positive root.


ANSWER.  W = 9 ft.  L = 12 ft.

Solved.